### v1

@@ 7.2 Synthetic Data
We generate three source distributions for a synthetic test. Each distribution is a 3-dimensional product of standard continuous distributions; see LABEL:sec-synthetic-data-setup for details, with full code available in this paper’s GitHub repository [Richardson_multiscale_paper_code]. Benchmarking the two approaches yields the following for the single-scale factorize function:
B
Range (min...max): 1.305 s...3.487 s   GC (min...max): 15.07%
@@
The function multiscale_factorize runs roughly seven times faster and uses about one-fourth of the memory. This comparison includes overhead such as interpolating the tensor and repeatedly calling the internal factorize function, which represents approximately $4\%$ of the total time. The specific numbers are less important than the qualitative comparison: both algorithms could be optimized and benchmarked on state-of-the-art hardware, but this basic implementation demonstrates that multiscale methods can accelerate single-scaled algorithm.
In these multiscale_factorize benchmarks, rather than performing one iteration at each scale coarser than the finest scale, we advance to the next scale once sufficient progress has been made; see LABEL:sec-synthetic-data-setup for details. This differs from our analysis in Corollary˜6.10, which suggests using exactly $K_{s}=1$ at coarse scales. However, running additional iterations at coarser scales often reduces the total number of iterations required at the finest scale.
Figure˜8 (left) shows median runtime as a function of the number of fixed coarse iterations $K_{s}$ in multiscale for $s=S,S-1,\dots,2$, where we run iterations at the finest scale until the objective falls below $10^{-6}$. This figure demonstrates that performing more than one iteration at coarser scales improves multiscale performance, but only to a point: in Fig.˜8 (left), more than 12 iterations at each coarse scale wastes time. Figure˜8 (right) illustrates this by showing the number of fine-scale iterations $K_{1}$ needed to converge to a minimum value. Our implementation of multiscale_factorize [Richardson_BlockTensorFactorization_jl] addresses this by allowing multiple iterations at coarser scales and advancing to the next scale when other criteria are met, such as small gradients or objective values.
Both plots in Fig.˜8 show that multiscale with exactly $K_{s}=1$ iterations at coarse scales $s=S,S-1,\dots,2$ performs slightly worse than single-scale optimization. We attribute this to interpolation and to overhead costs in the multiscale approach. The problem size of $2^{6}+1=65$ points along each continuous dimension may be too small to see immediate benefit from multiscale when only one iteration is performed at each scale; larger problems may needed to realize the advantage illustrated by Fig.˜1. However, flexible criteria to advance between scales may make multiscale competitive even for smaller problem.
@@ 7.3 Real Geological Data
We use the same sedimentary data and Tucker-$1$ model described in Graham et. al. [graham_tracing_2025] to factorize a tensor containing mixtures of estimated densities. LABEL:sec-geo-numerics-details provides the details, and the full code appears in distribution_unmixing_geology.jl in this paper’s GitHub repository [Richardson_multiscale_paper_code]. For factorize, we obtain
B
Range (min...max): 174.316 ms...593.727 ms   GC (min...max): 19.75%
@@
By every metric, the multiscale approach is faster and uses less memory than the standard single-scale factorization. Comparing median time and memory estimates, the multiscale method runs roughly four times as fast and uses about a third of the memory. These wall clock-times include significant garbage collection (GC), suggesting that design improvements such as preallocated arrays could reduce memory usage further. Such improvements would accelerate both factorize and multiscale_factorize. Nevertheless, we observe that multiscale_factorize triggers garbage collection less frequently because more computations occur on smaller arrays.
@@ 8 Conclusion
We developed a multiscale approach for optimizing Lipschitz continuous functions by solving discretized problems at progressively finer scales. Our convergence analysis establishes that both greedy and lazy variants achieve the finest-scale solution through per-scale descent combined with controlled interpolation error. The framework extends to linear and norm constraints and to tensor-valued functions.
@@
The explicit error bounds that we derive quantify the tradeoff between discretization granularity and solution accuracy, and prove that multiscale can outperform direct fine-scale optimization when the problem size and Lipschitz constant are sufficiently large.

### v2

@@ 7.2 Synthetic Data
We generate three source distributions for a synthetic test. Each distribution is a 3-dimensional product of standard distributions; see LABEL:sec-synthetic-data-setup this paper’s GitHub repository [21] for details. Benchmarking the two approaches across 20 samples yields the following for the single-scale factorize function:
R
@@
The function multiscale_factorize runs roughly seven times faster and uses about one-fourth of the memory. This comparison includes overhead such as interpolating the tensor and repeatedly calling the internal factorize function, which represents approximately $4\%$ of the total time. The specific numbers are less important than the qualitative comparison: both algorithms could be optimized and benchmarked on state-of-the-art hardware, but this basic implementation demonstrates that multiscale methods can accelerate single-scaled algorithms.
In these multiscale_factorize benchmarks, rather than performing one iteration at each scale coarser than the finest scale, we advance to the next scale once sufficient progress has been made; see LABEL:sec-synthetic-data-setup for details. This differs from our analysis in Corollary˜6.8, which suggests using exactly $K_{s}=1$ at coarse scales. However, running additional iterations at coarser scales often reduces the total number of iterations required at the finest scale.
Figure˜7 (left) shows median runtime as a function of the number of fixed coarse iterations $K_{s}$ in multiscale for $s=S,S-1,\dots,2$, where we run iterations at the finest scale until the objective falls below $10^{-6}$. This figure demonstrates that performing more than one iteration at coarser scales improves multiscale performance, but only to a point: in Fig.˜7 (left), more than 12 iterations at each coarse scale wastes time. Figure˜7 (right) illustrates this by showing the number of fine-scale iterations $K_{1}$ needed to converge to a minimum value. Our implementation of multiscale_factorize [20] addresses this by allowing multiple iterations at coarser scales and advancing to the next scale when other criteria are met, such as small gradients or objective values.
Both plots in Fig.˜7 show that multiscale with exactly $K_{s}=1$ iterations at coarse scales $s=S,S-1,\dots,2$ performs slightly worse than single-scale optimization. We attribute this to interpolation and to overhead costs in the multiscale approach. The problem size of $2^{6}+1=65$ points along each continuous dimension may be too small to see immediate benefit from multiscale when only one iteration is performed at each scale; larger problems may needed to realize the advantage illustrated by Fig.˜1. However, flexible criteria to advance between scales may make multiscale competitive even for smaller problems as shown in the benchmark tests.
@@ 7.3 Real Geological Data
We use the same sedimentary data and Tucker-$1$ model described in Graham et. al. [11] to factorize a tensor containing mixtures of estimated densities. LABEL:sec-geo-numerics-details provides the details, and the full code appears in distribution_unmixing_geology.jl in this paper’s GitHub repository [21]. For factorize and multiscale_factorize, we obtain the respective benchmarks:
R
@@
By every metric, the multiscale approach is faster and uses less memory than the standard single-scale factorization. Comparing median time and memory estimates, the multiscale method runs roughly four times faster with about a third of the memory.
@@ 8 Conclusion
We developed a multiscale approach for optimizing Lipschitz continuous functions by solving discretized problems at progressively finer scales. Our convergence analysis establishes that both greedy and lazy variants achieve the finest-scale solution through per-scale descent combined with controlled interpolation error. The framework extends to linear and norm constraints and tensor-valued functions.
@@
The explicit error bounds derived quantify the tradeoff between discretization granularity and solution accuracy, and prove that multiscale outperforms direct fine-scale optimization when the problem size and Lipschitz constant are sufficiently large.